OpenStudy (anonymous):

mkuma bought a pen drive and clock for $1300.He sold pen drive at the profit of 20% and clock at the profit of 25%.His net profit is (231/3)%.Find the cost price of clock.

OpenStudy (anonymous):

simultaneous equation problem Eq.1: pen + clock = 1300 express clock as f(pen) gives: clock = 1300 - pen Eq. 2: pen*1.2 + clock*1.25 = 1300*1.233 Substitute clock as a f(pen): pen*1.2 + (1300 - pen)*1.25 = 1300*1.233 1300*1.25 - 0.05pen = 1300*1.2333 1300*0.0167 = 0.05pen pen = 1300*0.0167)/(0.05) pen = 442 clock = 858

OpenStudy (anonymous):

perfect

OpenStudy (anonymous):

You can substitute the pen and clock values into Eq. 2 to double check :-)

OpenStudy (anonymous):

not to be judgemental, but is this really in the spirit of openstudy? if somone just get the answer here, their not learning anything

OpenStudy (anonymous):

Tried to respond with an answer to the specific question as well as an explanation of the method including checking :-)

OpenStudy (anonymous):

okay, i suppose thats good... but just for future reference, i imagine its better to solve a paralel example instead of the origonal problem